OpenStudy (anonymous):

Joey has a collection of baseball cards that he has separated into two different value categories. His common cards are worth $0.75 each and his premium cards are worth $2.00 each. Together, Joey’s entire baseball card collection of 88 cards is worth $106.00. How many cards in Joey’s collection are worth $0.75?

OpenStudy (mathstudent55):

Let x represent the number of $0.75 cards Let y represent the number of $2 cards Since he has 88 cards, then x + y = 88 Now we need an equation that shows the value of the cards. x number of $0.75 cards are worth 0.75x dollars y number of $2 cards are worth 2y dollars The entire collection is worth $106, so 0.75x + 2y = 106 Now we put these two equations together and we have a system of equations that we need to solve. x + y = 88 0.75x + 2y = 106 Do you know how to solve a system of equations in two variables?